OpenStudy (anonymous):

3x-2y+0 X+2y=-5 I'm confused. Its like the other one I posted.

OpenStudy (anonymous):

3x - 2y = 0 3x = 2y y = 1.5 x Place in eqtn 2 x + 3 x = -5 4x = -5 x = -5/4 But that is using substitution if that the answer LOLZ :D

OpenStudy (anonymous):

Multiply equation 2 by 3 to get 3x + 6y = -15, then take this away from eqequation 1. The x's are eliminated, so solve for y.

OpenStudy (anonymous):

Randi Candi, you're real smart. Thanksalot for helping me on this problem.

OpenStudy (anonymous):

you are VERY welcome :D :D Thank You XD XD